OpenStudy (anonymous):

Let f(x) = 4 + 5x and g(x) = 2x - 1. Find f(g(x)) and g(f(x)) Choices:) A.) f(g(x)) = 10x - 1; g(f(x)) = 10x + 7 B.) f(g(x)) = 7x + 3; g(f(x)) = 10x + 7 C.) f(g(x)) = -7x - 3; g(f(x)) = -10x + 7 D.) f(g(x)) = -10x - 7; g(f(x)) = 7x + 3

OpenStudy (anonymous):

f(g(x))=4+5(2x-1) distribute =4+10x-5 simplify =-1+10X g(f(x))=2(4+5x)-1 distribute =8+10x-1 simplify =7+10x

OpenStudy (anonymous):

So would the answer be A ?